# Graph Neural Networks in EEG-based Emotion Recognition: A Survey

### v5

@@ Abstract
Compared to other modalities, EEG-based emotion recognition can intuitively respond to the emotional patterns in the human brain and, therefore, has become one of the most concerning tasks in the brain-computer interfaces field. Since dependencies within brain regions are closely related to emotion, a significant trend is to develop Graph Neural Networks (GNNs) for EEG-based emotion recognition. However, brain region dependencies in emotional EEG have physiological bases that distinguish GNNs in this field from those in other time series fields. Besides, there is neither a comprehensive review nor guidance for constructing GNNs in EEG-based emotion recognition. In the survey, our categorization reveals the commonalities and differences of existing approaches under a unified framework of graph construction. We analyze and categorize methods from three stages in the framework to provide clear guidance on constructing GNNs in EEG-based emotion recognition. In addition, we discuss several open challenges and future directions, such as Temporal full-connected graph and Graph condensation.
@@ 1 Introduction
Emotion is an integral and complex aspect of human cognition that is decisive in human decision-making, behavior, and social interaction. Therefore, Emotion Recognition is essential in areas such as the diagnosis of mental disorders and Brain-Computer Interfaces (BCIs). Human emotions have various external manifestations, such as body language, voice, expression, and physiological signals. Among them, the Electroencephalogram (EEG) signals are perceived as a product of coordinated neural activity and electrical signal transmission within the human brain, which can directly and objectively reveal genuine human emotion compared to other manifestations and thus are known as one of the most unique and vital data in emotion recognition. EEG records emotional activity from different brain regions and, therefore, can reflect the complex interactions between brain regions in emotional states and the strong correlation between specific brain regions and specific emotions Min et al. (2022). In summary, inferring and exploiting complex dependencies between brain regions in emotional EEG has become a significant research direction in EEG-based emotion recognition.
Figure 1: Number of GNN-based methods and non-GNN-based methods, and their relative proportion in EEG-based emotion recognition in top journals and conferences over the last six years. So far, more than one-third of the methods utilize GNNs.
Graph Neural Networks (GNNs) emerge as a powerful tool for modeling dependencies of emotional EEG within the network neuroscience framework. As networks that manipulate graph-structured data, GNNs can effectively extract features utilizing dependencies between brain regions in emotional EEG. These dependencies represent the connectivity patterns and interactions of brain regions in emotional states, which are directly associated with specific emotional activities. Therefore, benefiting from its dependencies inference ability, GNNs designed for EEG-based emotion recognition can improve classification tasks compared to traditional analysis methods and potentially uncover new insights in neuroscience. As shown in Fig. 1, there has been a significant growth in the number and proportion of methods using GNNs in the EEG-based emotion recognition field over the last six years.
@@
  • We summarise and highlight future directions to facilitate GNN-based works in EEG-based emotion recognition.
@@ 2 Overview of Categorization
In recent years, EEG-based emotion recognition has seen an influx of methods with GNNs. These methods focus on different aspects of designing GNNs to infer dependencies within emotional EEG. They generally answer these three critical questions during the graph construction process: What features are chosen as nodes? How to calculate edges? Which graph structure is utilized? Therefore, categorizing these methods at the model level is not instructive nor sufficiently in-depth. To better understand these methods and propose a construction guideline for GNNs in EEG-based emotion recognition, our categorization is based on a unified framework with its corresponding stages.
@@ 2.1 Framework for specific GNNs
[…]
@@ 2.2 Stages of Framework
Node-level Stage indicates the process of selecting features as nodes V. Instead of focusing on the detailed composition and parameters of the node embedding extractor, this survey focuses on the composition of node features. Therefore, we categorize this stage according to the type of node features, as shown in Fig. 3 (a). These two categories are Univariate node, which utilizes a single feature, and Hybrid node, which employs multiple features. The details of each category will be explained in the Section. 3.
Edge-level Stage represents the calculation of the edge matrix E. The edge matrix represents the relationships between electrode nodes. Matrices representing such relationships differ in different methods, such as the adjacency matrix or the Laplace matrix. We uniformly regard them as edge matrices. Depending on whether the model parameters are involved in the calculation, we categorize the Edge-level Stage of the existing methods into two categories: Model-independent edge and Model-dependent edge, as shown in Fig. 3 (b). The sub-categories of each category will be explained in detail in Section. 4.
Graph-level Stage denotes the process of modeling different types of graph structures G. The basic units utilized by existing methods for reasoning about graph embedding are Graph Convolutional Networks (GCNs) or Graph Attention Networks (GATs). The perspective of this survey is not on the detailed model components but on the graph structure used to represent the dependencies within emotional EEG. The graph structure of existing methods can be categorized into four types, as shown in Fig 3 (c). Multi-graph indicates that the model employs a multi-stream structure to simultaneously construct different graphs to incorporate multiple dependencies of emotional EEG. Hierarchical graph denotes that the model organizes nodes into multiple groups to infer dependency between brain regions related to emotion. Time series graph is a series of graphs in the temporal dimension that represent the temporal dependency of emotional EEG. Sparse graph structure represents sparsely connected graphs that match the concentration of brain activity in emotional states. The sub-categories of each category will be explained in detail in Section. 5.
@@ 3.1 Univariate node
Temporal domain node implies that the model employs the temporal domain feature as nodes. This is the most commonly used node in the current methods, especially the differential entropy (DE) feature. The popularity of the DE feature, in addition to the fact that many open-source EEG-based emotion recognition datasets include this feature, such as SEED Zheng and Lu (2015), DE feature itself is suitable for measuring the uncertainty or randomness of the signal, which is adequate for describing the diverse dependencies in the emotional EEG. In addition to this, HD-GCN Ye et al. (2022) and SparseDGCNN Zhang et al. (2021) employ the Differential Asymmetry of Synchronisation for Multivariate Signals (DASM) feature.
@@
Raw Signal node represents directly using the raw emotional EEG as nodes. This node is the least common in EEG-based emotion recognition tasks because raw emotional EEG is highly individualized, which remains more emotion-independent disturbances. Besides, The dimension of Raw signal node is higher compared to other nodes. However, this node maximizes information retention and enables end-to-end training of the model.
@@
Node
Baseline
@@
LR-GCN Jin et al. (2021)
GECNN Song et al. (2021b)
TARDGCN Li et al. (2023a)
OMHGL Pan et al. (2023)
GIGN Ding and Guan (2023)
@@
HetEmotionNet Jia et al. (2021)
LGGNet Ding et al. (2023)
@@ 3.2 Hybrid node
Hybrid node implies that the node mixes a variety of features. It is consistent with the goal of Multi-graph mentioned in Section 5.1, which is inferring multiple dependencies within emotional EEG. The difference is that Multi-graph builds multiple graphs for different dependencies, while Hybrid nodes fuse features that contain different dependencies. Hybrid nodes can utilize the complementarity between different information to provide a more comprehensive view of emotional EEG dependencies. However, it increases the dimension of the node. Besides, due to the different distributions of features, it can easily lead to overfitting when the amount of data is insufficient. Hybrid node contains two categorizations. CR-GAT Liu et al. (2022b), DBGC-ATFFNet-AFTL Sun et al. (2022), and EGFG Li et al. (2022a) employ Temporal&Frequency node that fuses DE features, and PSD features to infer both temporal and frequency dependencies simultaneously. Methods like ERHGCN Zheng et al. (2021) utilize Aggregation node that aggregates a variety of features, including DASM, Average Size of Messages (ASM), etc., to preserve as many dependencies as possible.
@@ 4.1 Model-independent edge
Model-independent edge implies that the calculation of the edge does not include the model parameters and backpropagation does not update this edge. It is based on existing brain physiology paradigms or signal processing paradigms. The fact that such edges are independent of the data distribution and model and conform to physiological or signaling patterns increases the plausibility and stability of Model-independent edge. Moreover, such an edge does not increase model parameters and is compatible with any model structure. The Model-independent edge contains three edges, as shown in Table. 2
Prior edge refers to the edge that is set manually based on prior knowledge. Prior Knowledge refers to existing physiological paradigms about electrode connections, which is physiologically plausible. However, emotional EEG has individual differences, so the connections between electrodes differ between subjects. This makes the rationality of Prior edge limited because it cannot be automatically updated. Based on prior knowledge that the strength of connections between brain regions decays as a function of the inverse square of the physical distance Salvador et al. (2005), RGNN Zhong et al. (2020), and DAGAM Xu et al. (2023) use 3D co-ordinates to calculate the physical distances between electrodes as the edges. The International 10-20 EEG Electrode Standard depicts the connectivity between electrodes, and MD-GCN Du et al. (2022) and SFE-Net Deng et al. (2021) transform it into an edge matrix. LGGNet Ding et al. (2023) limits the connectivity between electrodes to specific brain regions based on the distribution of brain regions.
Distance edge refers to treating signals as vectors and taking vector distances as edges. Calculating the distance between vectors is commonly used to measure the relationship between signals. Distance edge can be updated as the node changes, but this relationship is physiologically independent of emotion. TARDGCN Li et al. (2023a) and OMHGL Pan et al. (2023) apply Cosine distance between nodes as edges. Similarly, EGFG Li et al. (2022a) and JSCFE Peng et al. (2022b) utilize Euclidean distance. MST-GNN Liu et al. (2022a) uses the Phase Lag Index, which analyses the phase distance between node signals.
Similarity edge refers to calculating the similarity of signals as edges. It differs from Distance edge only in the mathematical meaning, so Similarity edge is dynamic but does not always conform to the physiological paradigm. MD-AGCN Li et al. (2021b) and ST-GCLSTM Feng et al. (2022) utilize the Pearson Correlation Coefficient to compute the strength of the linear relationship between signals as their edge. HB-SR Yang et al. (2024) uses Phase Locking Value to calculate the degree of phase similarity between electrode signals at a particular frequency as their edge. In addition, HD-GCN Ye et al. (2022) applies the Gaussian similarity, and HetEmotionNet Jia et al. (2021) indicates mutual information to represent the correlation between two signals as the edge.
@@
Edge
Baseline
@@
RGNN Zhong et al. (2020)
SFE-Net Deng et al. (2021)
MD-GCN Du et al. (2022)
LGGNet Ding et al. (2023)
DAGAM Xu et al. (2023)
@@
EGFG Li et al. (2022a)
JSCFE Peng et al. (2022b)
MST-GNN Liu et al. (2022a)
TARDGCN Li et al. (2023a)
OMHGL Pan et al. (2023)
@@
MD-AGCN Li et al. (2021b)
HetEmotionNet Jia et al. (2021)
ST-GCLSTM Feng et al. (2022)
HD-GCNYe et al. (2022)
HB-SRYang et al. (2024)
@@ 4.2 Model-dependent edge
Model-dependent edge refers to the computation of edges using model parameters. The edges containing model parameters can be dynamically updated with training so that the model can be generalized over different data. Compared to Model-independent edges that are limited by prior knowledge or vector relationships, Model-dependent edge can capture more complex brain region relationships in emotional EEG without manually adjusting when handling different signals. However, additional parameters make it challenging to balance accuracy and computational burden. Model-dependent edge contains four categories, as shown in Table. 3.
@@
where $\rho$ denotes the learning rate. $\frac{\partial\mathcal{L}}{\partial\mathbf{E}}$ refers to the partial derivative of the loss $\mathcal{L}$ to $\mathbf{E}_{i,j}$.
@@
where $\mathbf{W}$ represents the model parameter utilized as weight. $\xi(\cdot)$ is activation function. $f[\cdot,\cdot]$ denotes the relationship function between nodes. Siam-GCAN Zeng et al. (2022) and DBGC-AFFNet-AFTL Sun et al. (2022) utilize distance between nodes as the relationship function. LR-GCN Jin et al. (2021) and STFCGAT Li et al. (2023b) concatenate any two nodes to perform a linear transformation as the relationship function.
@@
where $\mathbf{W}$ represents the model parameter utilized as bias. MRGCN Qiu et al. (2023) utilizes the calculation of the physical distance between nodes as the relationship function. OGSSL Peng et al. (2022a) applies the similarity between nodes as the relationship function.
Subspace edge denotes the projection of nodes to subspace by parameters to obtain similarity as edges. It is similar in principle to the Attention mechanism and allows the model to capture the complex relationship between every two nodes. However, it adds more model complexity and computational cost than other Model-dependent edges. MD-AGCN Li et al. (2021b), JSCFE Peng et al. (2022b), and ASTG-LSTM Li et al. (2021c) use the following to implement Subspace edge:
@@
Edge
Baseline
@@
DGCNN Song et al. (2018)
SGA-LSTM Liu et al. (2019)
GCB-Net Zhang et al. (2019)
HD-GCN Ye et al. (2022)
AHGCN Xue et al. (2022)
GIGN Ding and Guan (2023)
EmoGT Jiang et al. (2023)
@@
LR-GCN Jin et al. (2021)
Siam-GCAN Zeng et al. (2022)
DBGC-ATFFNet-AFTL Sun et al. (2022)
STFCGAT Li et al. (2023b)
@@
OGSSL Peng et al. (2022a)
MRGCN Qiu et al. (2023)
@@
MD-AGCN Li et al. (2021b)
ASTG-LSTM Li et al. (2021c)
JSCFE Peng et al. (2022b)

### v6

@@ Abstract
Compared to other modalities, EEG-based emotion recognition can intuitively respond to the emotional patterns in the human brain and, therefore, has become one of the most concerning tasks in the brain-computer interfaces. Since dependencies within brain regions are closely related to emotion, a significant trend is to develop Graph Neural Networks (GNNs) for EEG-based emotion recognition. However, brain region dependencies in emotional EEG have physiological bases that distinguish GNNs in this field from those in other time series fields. Besides, there is neither a comprehensive review nor guidance for constructing GNNs in EEG-based emotion recognition. In the survey, our categorization reveals the commonalities and differences of existing approaches under a unified framework of graph construction. We analyze and categorize methods from three stages in the framework to provide clear guidance on constructing GNNs in EEG-based emotion recognition. In addition, we discuss several open challenges and future directions, such as temporal fully-connected graph and graph condensation.
@@ 1 Introduction
Emotion is an integral and complex aspect of human cognition that is decisive in human decision-making, behavior, and social interaction. Therefore, Emotion Recognition is essential in areas such as the diagnosis of mental disorders and Brain-Computer Interfaces (BCIs). Human emotions have various external manifestations, such as body language, voice, expression, and physiological signals. Among them, the Electroencephalogram (EEG) signals are perceived as a product of coordinated neural activity and electrical signal transmission within the human brain, which can directly and objectively reveal genuine human emotion compared to other manifestations and thus are known as one of the most unique and vital data in emotion recognition. EEG records emotional activity from different brain regions and, therefore, can reflect the complex interactions between brain regions in emotional states and the strong correlation between specific brain regions and specific emotions [25]. In summary, inferring and exploiting complex dependencies between brain regions in emotional EEG has become a significant research direction in EEG-based emotion recognition.
Graph Neural Networks (GNNs) emerge as a powerful tool for modeling dependencies of emotional EEG within the network neuroscience framework. As networks that manipulate graph-structured data, GNNs can effectively extract features utilizing dependencies between brain regions in emotional EEG. These dependencies represent the connectivity patterns and interactions of brain regions in emotional states, which are directly associated with specific emotional activities. Therefore, benefiting from their dependencies inference ability, GNNs designed for EEG-based emotion recognition can improve classification tasks compared to traditional analysis methods and potentially uncover new insights in neuroscience. There has been a significant growth in the number and proportion of methods using GNNs in the EEG-based emotion recognition field over the last six years.
@@
  • We summarize and highlight future directions to facilitate GNN-based works in EEG-based emotion recognition.
@@ 2 Overview of Categorization
In recent years, EEG-based emotion recognition has seen an influx of methods with GNNs. These methods focus on different aspects of designing GNNs to infer dependencies within emotional EEG. They generally answer these three critical questions during the graph construction process: What features are chosen as nodes? How to calculate edges? Which graph structure is utilized? Therefore, categorizing these methods at the model level is neither instructive nor sufficiently in-depth. To better understand these methods and propose a construction guideline for GNNs in EEG-based emotion recognition, our categorization is based on a unified framework with its corresponding stages.
@@ 2.1 Framework for specific GNNs
EEG-based emotion recognition task for GNNs can be indicated as taking emotional EEG $X\in\mathbb{R}^{C\times S}$ and corresponding label $Y\in\mathbb{R}^{1}$ as input, constructing graphs and predicting emotional labels $Y^{\prime}\in\mathbb{R}^{1}$ via graph embeddings. Let $\textbf{G}(\textbf{V},\textbf{E})$ denote a graph. $\textbf{V}\in\mathbb{R}^{N_{v}\times D}$ represents the node set, where $N_{v}=C$ indicates the number of channels and $D$ denotes node embedding dimension. $\textbf{E}\in\mathbb{R}^{N_{e}}$ represents the edge matrix, where $N_{e}=C^{2}$ denotes the number of edges. The goal of the model can be represented as learning a function $f(\textbf{G})\rightarrow Y^{\prime}$ that maps the graph to the corresponding emotional label. Therefore, the unified framework for building GNNs in EEG-based emotion recognition consists of three stages, as shown in Fig. 1. First, the Node-level Stage represents the model selection of features to be used as nodes V. Second, the Edge-level Stage indicates that the model computes the edge matrix E. Finally, the Graph-level Stage denotes the process of graph G construction.
@@ 2.2 Stages of Framework
Node-level Stage indicates the process of selecting features as nodes V. Instead of focusing on the detailed composition and parameters of the node embedding extractor, this survey focuses on the composition of node features. Therefore, we categorize this stage according to the type of node features, as shown in Fig. 2 (a). These two categories are Univariate node, which utilizes a single feature, and Hybrid node, which employs multiple features. The details of each category will be explained in Section 3.
Edge-level Stage represents the calculation of the edge matrix E. The edge matrix represents the relationships between electrode nodes. Matrices representing such relationships differ in different methods, such as the adjacency matrix or the Laplacian matrix. We uniformly regard them as edge matrices. Depending on whether the model parameters are involved in the calculation, we categorize the Edge-level Stage of the existing methods into two categories: Model-independent edge and Model-dependent edge, as shown in Fig. 2 (b). The sub-categories of each category will be explained in detail in Section 4.
Graph-level Stage denotes the process of modeling different types of graph structures G. The basic units utilized by existing methods for reasoning about graph embedding are Graph Convolutional Networks (GCNs) or Graph Attention Networks (GATs). The perspective of this survey is not on the detailed model components but on the graph structure used to represent the dependencies within emotional EEG. The graph structure of existing methods can be categorized into four types, as shown in Fig. 2 (c). Multi-graph indicates that the model employs a multi-stream structure to simultaneously construct different graphs to incorporate multiple dependencies of emotional EEG. Hierarchical graph denotes that the model organizes nodes into multiple groups to infer dependency between brain regions related to emotion. Time series graph is a series of graphs in the temporal dimension that represent the temporal dependency of emotional EEG. Sparse graph structure represents sparsely connected graphs that match the concentration of brain activity in emotional states. The sub-categories of each category will be explained in detail in Section 5.
@@ 3.1 Univariate node
Temporal domain node implies that the model employs the temporal domain feature as nodes. This is the most commonly used node in the current methods, especially the differential entropy (DE) feature. The popularity of the DE feature, in addition to the fact that many open-source EEG-based emotion recognition datasets include this feature, such as SEED [46], DE feature itself is suitable for measuring the uncertainty or randomness of the signal, which is adequate for describing the diverse dependencies in the emotional EEG. In addition to this, HD-GCN [38] and SparseDGCNN [40] employ the Differential Asymmetry of Synchronisation for Multivariate Signals (DASM) feature.
@@
Raw Signal node represents directly using the raw emotional EEG as nodes. This node is the least common in EEG-based emotion recognition tasks because raw emotional EEG is highly individualized, which retains more emotion-independent disturbances. Besides, the dimension of Raw signal node is higher compared to other nodes. However, this node maximizes information retention and enables end-to-end training of the model.
@@
Node Baseline Node Feature
@@
LR-GCN [10]
GECNN [32]
TARDGCN [14]
OMHGL [26]
GIGN [3]
EmT [5]
Power Spectral Density
Hilbert-Huang transform spectrum
Welch
Power Spectral Density
Power Spectral Density
Power Spectral Density
@@
HetEmotionNet [8]
LGGNet [4]
RPGCN [48]
VSGT [20]
VBH-GNN [21]
Time-domain Amplitude
Power Conversion
Linear Embedding
Time-domain Amplitude
Time-domain Amplitude
@@ 3.2 Hybrid node
Hybrid node implies that the node mixes a variety of features. It is consistent with the goal of Multi-graph mentioned in Section 5.1, which is inferring multiple dependencies within emotional EEG. The difference is that Multi-graph builds multiple graphs for different dependencies, while Hybrid nodes fuse features that contain different dependencies. Hybrid nodes can utilize the complementarity between different information to provide a more comprehensive view of emotional EEG dependencies. However, it increases the dimension of the node. Besides, due to the different distributions of features, it can easily lead to overfitting when the amount of data is insufficient. Hybrid node contains two categorizations. CR-GAT [23], DBGC-ATFFNet-AFTL [34], and EGFG [11] employ Temporal&Frequency node that fuses DE features, and PSD features to infer both temporal and frequency dependencies simultaneously. Methods like ERHGCN [45] utilize Aggregation node that aggregates a variety of features, including DASM, Average Size of Messages (ASM), etc., to preserve as many dependencies as possible.
@@ 4.1 Model-independent edge
Model-independent edge implies that the calculation of the edge does not include the model parameters and backpropagation does not update this edge. It is based on existing brain physiology paradigms or signal processing paradigms. The fact that such edges are independent of the data distribution and model and conform to physiological or signaling patterns increases the plausibility and stability of Model-independent edge. Moreover, such an edge does not increase model parameters and is compatible with any model structure. The Model-independent edge contains three edges, as shown in Table 2
Prior edge refers to the edge that is set manually based on prior knowledge. Prior Knowledge refers to existing physiological paradigms about electrode connections, which is physiologically plausible. However, emotional EEG has individual differences, so the connections between electrodes differ between subjects. This makes the rationality of Prior edge limited because it cannot be automatically updated. Based on prior knowledge that the strength of connections between brain regions decays as a function of the inverse square of the physical distance [30], RGNN [47], and DAGAM [35] use 3D co-ordinates to calculate the physical distances between electrodes as the edges. The International 10-20 EEG Electrode Standard depicts the connectivity between electrodes, and MD-GCN [6] and SFE-Net [2] transform it into an edge matrix. LGGNet [4] limits the connectivity between electrodes in specific brain regions based on the distribution of brain regions.
Distance edge refers to treating signals as vectors and taking vector distances as edges. Calculating the distance between vectors is commonly used to measure the relationship between signals. Distance edge can be updated as the node changes, but this relationship is physiologically independent of emotion. TARDGCN [14] and OMHGL [26] apply Cosine distance between nodes as edges. Similarly, EGFG [11] and JSCFE [28] utilize Euclidean distance. MST-GNN [22] uses the Phase Lag Index, which analyses the phase distance between node signals.
Similarity edge refers to calculating the similarity of signals as edges. It differs from Distance edge only in the mathematical meaning, so Similarity edge is dynamic but does not always conform to the physiological paradigm. MD-AGCN [13] and ST-GCLSTM [7] utilize the Pearson Correlation Coefficient to compute the strength of the linear relationship between signals as their edge. HB-SR [37] uses Phase Locking Value to calculate the degree of phase similarity between electrode signals at a particular frequency as their edge. In addition, HD-GCN [38] applies the Gaussian similarity, and HetEmotionNet [8] indicates mutual information to represent the correlation between two signals as the edge.
@@
Edge Baseline Edge Feature
@@
RGNN [47]
SFE-Net [2]
MD-GCN [6]
LGGNet [4]
DAGAM [35]
Physical Distances
Connectivity through Edge Matrix
Connectivity through Edge Matrix
Brain Region Connectivity
Physical Distances
@@
EGFG [11]
JSCFE [28]
MST-GNN [22]
TARDGCN [14]
OMHGL [26]
Euclidean distance
Euclidean distance
Phase Lag Index
Cosine distance
Cosine distance
@@
MD-AGCN [13]
HetEmotionNet [8]
ST-GCLSTM [7]
HD-GCN [38]
HB-SR [37]
Pearson Correlation Coefficient
Mutual Information
Pearson Correlation Coefficient
Gaussian Similarity
Phase Locking Value
@@ 4.2 Model-dependent edge
Model-dependent edge refers to the computation of edges using model parameters. The edges containing model parameters can be dynamically updated with training so that the model can be generalized over different data. Compared to Model-independent edges that are limited by prior knowledge or vector relationships, Model-dependent edge can capture more complex brain region relationships in emotional EEG without manually adjusting when handling different signals. However, additional parameters make it challenging to balance accuracy and computational burden. Model-dependent edge contains four categories, as shown in Table 3.
@@
where $\rho$ denotes the learning rate. $\frac{\partial\mathcal{L}}{\partial\mathbf{E}}$ refers to the partial derivative of the loss $\mathcal{L}$ with respect to $\mathbf{E}_{i,j}$.
@@
where $\mathbf{W}$ represents the model parameter utilized as weight. $\xi(\cdot)$ is activation function. $f[\cdot,\cdot]$ denotes the relationship function between nodes. Siam-GCAN [39] and DBGC-ATFFNet-AFTL [34] utilize distance between nodes as the relationship function. LR-GCN [10] and STFCGAT [19] concatenate any two nodes to perform a linear transformation as the relationship function.
@@
where $\mathbf{W}$ represents the model parameter utilized as bias. MRGCN [29] utilizes the calculation of the physical distance between nodes as the relationship function. OGSSL [27] applies the similarity between nodes as the relationship function.
Subspace edge denotes the projection of nodes to subspace by parameters to obtain similarity as edges. It is similar in principle to the Attention mechanism and allows the model to capture the complex relationship between every two nodes. However, it adds more model complexity and computational cost than other Model-dependent edges. MD-AGCN [13], JSCFE [28], HGCN [44], VBH-GNN [21] and ASTG-LSTM [15] use the following to implement Subspace edge:
@@
Edge Baseline Edge Feature
@@
DGCNN [33]
GCB-Net [42]
HD-GCN [38]
AHGCN [36]
GIGN [3]
EmoGT [9]
EmT [5]
DTC-GCN [49]
RPGCN [48]
VSGT [20]
Gaussian similarity
Learned Topology
Gaussian similarity
Learned Topology
Learned Topology
Learned Topology
Multi-view Learning
Sparse Causality
Relational Probabilistic
Prior Constrained Dependencies
@@
Siam-GCAN [39]
DBGC-ATFFNet-AFTL [34]
STFCGAT [19]
LR-GCN [10]
Distance-based Weights
Distance-based Weights
Linear Transformation
Attention Weights & Linear Transformation
@@
OGSSL [27]
MRGCN [29]
Euclidean Similarity
Physical Distance
@@
MD-AGCN [13]
ASTG-LSTM [15]
HGCN [44]
VBH-GNN [21]
Pearson Correlation Coefficient
Attention Mechanism
Heterogeneous Multi-head
Variational Bayesian Graph Inference